The magnetic field at the centre of a circular loop of wire of radius 1.43 cm is 3.41 x 104 T in magnitude. (a) What is the current in the wire? (b) If the loop is in a horizontal plane and the field at its centre is upward, is the current flowing clockwise, or counterclockwise, as viewed from above? clockwise counterclockwise
Answer #1

r = 1.43 cm

B = 3.41*10^-4 T

(a) As B = uo*I/(2R) ======> I = 2RB/uo = (2*0.0143*3.41*10^-4)/(4pi*10^-7) = 7.76 A

(b) Counterclockwise